OpenStudy (anonymous):

At 1.00 atm and 0 °C, a 5.04 L mixture of methane (CH4) and propane (C3H8) was burned, producing 16.5 g of CO2. What was the mole fraction of each gas in the mixture? Assume complete combustion.

OpenStudy (abb0t):

Keep in mind that combustion is defined by: C\(\sf \color{orange }{_n}\)H\(\sf \color{orange }{_{2n+2}}\) + O\(_2\) \(\rightarrow\) H\(_2\)O + CO\(_2\) where \(\sf \color{red}{n}\)= number of carbon(s)! However, notice that for your problem, you have TWO organics. Methane and Propane! Next, assume that the gases are ideal, so that you can use the formular \(\sf \color{blue}{PV = nRT}\), where P = pressure [atm] V = volume [L] n = mol R = gas constant [0.0821 \(\frac{liter~ \times ~atm}{mol ~ \times~K}\)] T = temperature [K] Now, from algebra, rearrange to solve for \(\sf \color{purple}{n}\)!! Which means you divide presure and volume by the ideal gas constant and temperature. (YOU'RE DOING THI FOR \(\sf \color{red}{BOTH}\) propane \(and\) methane [NOTE #2: or you you could use the fact that an ideal gas at STP = 22.4 \(\frac{L}{mol}\) Eiehter method, you get same amount of mol value. Now, this will get tricky, you have: \(\sf \color{green}a\)CH\(\sf \color{orange }{_{4}}\) + \(\sf \color{purple}b\)C\(\sf \color{red }{_3}\)H\(\sf \color{red}{_{8}}\) + O\(_2\) \[\left\{\begin{matrix} a+b = n & & \\ 4a + 8b =n& & \end{matrix}\right.\] the TOP = # of carbon on reactant side BOTTOM = # hydrogen on reactant side. and n = mol that you calculated above either using ideal gas law formula OR ideal volume @ STP. fROM ALGEBRA, you should know how to solve for two unknowns. You can use \(\sf \color{red}{substitution}\) method or \(\sf \color{green}{elimination}\) method. Finally, to get the mole fraction IS mol of methane divide by total mol. And you're done.